# CivRealm: A Learning and Reasoning Odyssey in Civilization for Decision-Making Agents

### 19 Jan 2024 v1

@@ 1 Introduction
In this paper, we present CivRealm, an interactive environment designed to push the boundaries of decision-making agents’ learning and reasoning. CivRealm is inspired by the iconic game Civilization, where each agent acts as a player to guide a civilization, mirroring the course of human history. Within this game, players are confronted with decisions including resource management, strategic planning, diplomacy, and warfare. The game rules profoundly align with the mechanics governing human society, and decision-making in the game span from long-term strategic vision (e.g., development prioritization) to fine-grained tactical maneuvers (e.g., unit control) (Figure 1).
@@ 4.2 BaseLang: Baseline Language-based Agent
Necessity and challenges The need for developing an LLM-based agent in CivRealm arises from two factors. 1) LLMs are capable of task generation and problem-solving, augmented by its extensive human knowledge base. Their capabilities and knowledge are promising for solving decision-making problems. 2) LLM is not yet sufficiently sophisticated for complex reasoning [77; 7]. In CivRealm, LLM is advantageous due to its ability to use natural language, prioritize strategic gameplay, and handle diplomatic interactions. However, constructing such an agent faces several challenges, including managing multiple in-game roles, handling sparse and complex observations, addressing the long-term impact of actions, and improving the agent’s decision-making over time.
@@ 5.1 Tensor-based Reinforcement Learning
To evaluate the performance of our tensor baseline across various mini tasks and full games, we kept the same architecture and training hyper-parameters and trained our models from scratch.
@@ 6 Conclusion
We present CivRealm, a distinctive challenge for decision-making agents, placing simultaneous emphasis on learning and reasoning abilities, which are evaluated across various mini-games and full games using diverse metrics. Experiments in CivRealm unveils limitations of contemporary methods, with tensor methods often yielding myopic strategies, and LLM-based agents still struggling with intricate reasoning tasks. To surmount these challenges, future research directions could involve integrating the strengths of RL and LLM, creating decision-making agents that combine adaptability and sophisticated reasoning. CivRealm serves as an ideal testing ground for these approaches.
Ackowledgement
This work was supported by the National Natural Science Foundation of China. The work of CivRealm is built upon several open-source projects, including Freeciv [74], Freeciv-web [75], FCIV-NET [73], and Freeciv-bot [57].
@@
The benchmark is committed to fostering cultural and social inclusivity. It is developed upon the open-source game Freeciv [74], with a conscious effort made to incorporate a wide array of civilizations. Should any instances of marginalization or misrepresentation arise, we encourage and facilitate contributions to the source code for rectification.
@@ A.1 More on Full Game and CivRealm Features
Dynamic space. As the game unfolds, players continuously produce new units, construct additional cities, engage in battles, and conquer other players. Consequently, the state and action space of a player undergo dynamic changes throughout the gameplay. Designing an effective decision model for the agent to adapt to this evolving space presents a significant challenge.
@@ A.1.3 Evaluation Metrics
CivRealm adopts game scores across 16 dimensions provided by the game engine to assess playing performance, including population, economics, production, cities, researched technologies, military units, wonders, research speed, land area, settled area, gold, units built, units killed, units lost, and units used, as well as an aggregated score for the overall performance evaluation. This aggregated score aggregates multiple factors, i.e., population, researched technologies, wonders, units built and killed, culture, and if spaceship is achieved by the player.
@@ A.2 Mini-game Details
We provide the concrete definition of each mini-game in Table 8. The mini-games belong to three classes -development, battle and diplomacy, covering a diversity of typical sub-tasks of the Civilization game. The development class of mini-games are closely related to city tile’s combined resource outputs (food, production, and trade), while the battle class of mini-games focus more on unit movement/attack operations. The player should be aware of the unit attack/defense/HP points to make full use of them. As for the last class of mini-game: diplomacy, it is about the art of trade and negotiation to achieve maximal interests for the player’s civilization.
@@ A.2.1 Victory Conditions, Score/Reward Settings and Difficulty Levels
Based on each mini-game’s characteristic, we design a set of evaluation principles for them. For instance, we assess development mini-games by verifying whether the combined resource outputs (food, production, and trade) from city tiles surpass a predetermined threshold. The success of battle mini-games is determined by the successful annihilation of enemy units or the conquest/defense of designated cities. We determine the achievement of diplomacy mini-games based on whether the agent negotiates favorable agreements with other players, such as exchanging valuable technologies. For the details of reward settings and the computation of victory conditions, please refer to Table 9.
We additionally partition the mini-games into different difficulty levels (easy, normal and hard) to better evaluate the models’ capability as well as enable curriculum design in future work. Specific criteria for the difficulty level divisions can be found in Table 9.
@@ A.2.2 Mini-game Randomization
To ensure sufficiently diverse and balanced mini-game instances, we conduct a set of randomization processing during mini-game generation. Table 10 gives exact perspectives on which randomization operations are performed for each type of mini-games.
@@
Table 9: Score, victory, reward and difficulty of mini-games.
@@ B.1 Observation
Tensor environment provides API for tensor models. Observation of tensor environment is a big tensor including all observations of map, unit, city, diplomacy, government, and technology in § A.1.1.
@@ B.2 Action
There are five types of actor in tensor environment, i.e., unit, city, diplomacy, government, and technology. Each type of actor has its corresponding action space and the overall action space is the joint action space of all actor types. For an actor, each action in its action space is denoted by an action class name associated with parameters. For example, an action in action class Go to a target tile is denoted by the action key: goto_num where goto is the action class name and num is a parameter describing the direction, thus action Go to the north tile is denoted as goto_1 where the parameter 1 corresponds to the north direction.
A chosen action returned to tensor environment is a triplet including the actor type, actor ID, and the action key. For example, assuming action Go to the north tile is chosen for a unit whose ID is 121, then the action triplet returned to tensor environment is (unit, 121, goto_1).
@@ C.1 Observation
Please refer to Figure 10 as an example of an actor’s observation dictionary. Zoomed-in observation corresponds to a $5\times 5$ mini-map centered on the actor’s location, thus it contains 25 tiles. Zoomed-out observation corresponds to a $15\times 15$ sub-map. In this example, we define that a block is made up of $5\times 5$ tiles, thus the above $5\times 5$ mini-map is the block where the actor is currently located, and the remaining tiles in zoomed-out observation make up the 8 surrounding blocks. Based on these settings, we get the observation of the actor. Each tile or block has a list containing its status, terrain, and the infrastructures on it. Besides, the information of units and cities, as well as their owners is also on the list. Unit and city owners are also associated with diplomacy tags, for example, units on current_tile belong to myself player; 3 cities in block_north_1 belong to an Alliance player.
@@ C.2 Action
The action returned to language environment is a triplet including the type of the actor, actor ID, and the name of the chosen action in available_actions. For example, in Figure 10, the type of the actor is unit, its actor ID is 121, assuming that action Go to the north tile has been chosen, then the final action return to language environment is (unit, 121, move North).
@@ Appendix D Estimation of State / Action Space Sizes
On Turn 5. Under ruleset Civ2Civ3, each nation started with 2 Settlers, 2 Workers and an Explorer. On turn 5, each Settler has acted at most 5 times, so it could appear at 121 possible places as itself, and could build a city at 81 positions. Each city originally had no less than 5 building options, and 24 working options. Each Worker, at the same time, could occur at similar positions, and when some work is done, the range it occurs will be smaller. Explorers could move 3 tiles per turn. So state space for 2 Settlers (or the cities they turned into) is about $\frac{1}{2}((121+81*5*24)^{2}-1521*5*24-121)\approx 5\times 10^{7}$ large (square of single Settlers status, counted in cases: a city is not built / a city is built, with city options, then excluding all conflicting cases where cities are too close to each other. Halving means two Settlers are identical in practice). Workers has states about $\frac{1}{2}((121+81*3)^{2}-121-81*3)=66066$. Explorer contributes $96131^{2}=$ 961 states. So total state space is of size about $3\times 10^{15}$.
@@ E.2.1 Challenges
Sparse and complex observation. In the early stages of the game, there are extensive unseen areas in the player’s field of view. This makes it necessary for early-stage agents to first learn to efficiently explore the environment using powerful prior knowledge, posing a significant challenge in constructing these agents. Additionally, the vast map size (4000-tile maps for our full game setup) and complex information on each tile overwhelms the context window of all language models. Thus, LLM agents have to prioritize reading the map properly and efficiently.
@@ E.2.2 Structure
We adopted an AutoGPT-like approach to build a baseline language-based agent. AutoGPT [72] is a memory-equipped automated LLM agent capable of making autonomous decisions in-context. The following sections will delve into observation, reasoning, commands, and structure separately. ion.

### 12 Mar 2024 v2

@@ 1 Introduction
In this paper, we present CivRealm, an interactive environment designed to push the boundaries of decision-making agents’ learning and reasoning. CivRealm is inspired by the iconic game Civilization, where each agent acts as a player to guide a civilization, mirroring the course of human history. Within this game, players are confronted with decisions including resource management, strategic planning, diplomacy, and warfare. The game rules profoundly align with the mechanics governing human society, and decision-making in the game span from long-term strategic vision (e.g., development prioritization) to fine-grained tactical maneuvers (e.g., unit control) (Figure 1). $*{}^{*}$qual contribution. ${}^{\dagger}$ Corresponding author.
@@ 4.2 BaseLang: Baseline Language-based Agent
Necessity and challenges The need for developing an LLM-based agent in CivRealm arises from two factors. 1) LLMs are capable of task generation and problem-solving, augmented by their extensive human knowledge base. Their capabilities and knowledge are promising for solving decision-making problems. 2) LLM is not yet sufficiently sophisticated for complex reasoning [77; 7]. In CivRealm, LLM is advantageous due to its ability to use natural language, prioritize strategic gameplay, and handle diplomatic interactions. However, constructing such an agent faces several challenges, including managing multiple in-game roles, handling sparse and complex observations, addressing the long-term impact of actions, and improving the agent’s decision-making over time.
@@ 5.1 Tensor-based Reinforcement Learning
To evaluate the performance of our tensor baseline across various mini-tasks and full games, we kept the same architecture and training hyper-parameters and trained our models from scratch.
@@ 6 Conclusion
We present CivRealm, a distinctive challenge for decision-making agents, placing simultaneous emphasis on learning and reasoning abilities, which are evaluated across various mini-games and full games using diverse metrics. Experiments in CivRealm unveil limitations of contemporary methods, with tensor methods often yielding myopic strategies, and LLM-based agents still struggling with intricate reasoning tasks. To surmount these challenges, future research directions could involve integrating the strengths of RL and LLM, creating decision-making agents that combine adaptability and sophisticated reasoning. CivRealm serves as an ideal testing ground for these approaches.
Acknowledgment
The work of CivRealm is built upon several open-source projects, including Freeciv [74], Freeciv-web [75], FCIV-NET [73], and Freeciv-bot [57].
This work was supported by the National Natural Science Foundation of China. This work was also supported by Wuhan East Lake High-Tech Development Zone, National Comprehensive Experimental Base for Governance of Intelligent Society.
@@ Acknowledgment
CivRealm can potentially contribute to the development of more socially aware AI agents, which could have applications in areas like policy-making, economic simulations, conflict resolution, and education. Moreover, the learning from such simulations could offer insights into the dynamics of human society, historical event outcomes, and potential future societal trajectories. Overall, the societal impacts of CivRealm and similar social simulation platforms can be profound, aiding in the development of AI that is more aligned with human values and societal goals.
The development of CivRealm is committed to fostering cultural and social inclusivity. It is developed upon the open-source game Freeciv [74], with a conscious effort made to incorporate a wide array of civilizations. Should any instances of marginalization or misrepresentation arise, we encourage and facilitate contributions to the source code for rectification.
@@ A.1 More on Full Game and CivRealm Features
Dynamic space. As the game unfolds, players continuously produce new units, construct additional cities, engage in battles, and conquer other players. Consequently, the state and action space of a player undergoes dynamic changes throughout the gameplay. Designing an effective decision model for the agent to adapt to this evolving space presents a significant challenge.
@@ A.1.3 Evaluation Metrics
CivRealm adopts game scores across 16 dimensions provided by the game engine to assess playing performance, including population, economics, production, cities, researched technologies, military units, wonders, research speed, land area, settled area, gold, units built, units killed, units lost, and units used, as well as an aggregated score for the overall performance evaluation. This aggregated score aggregates multiple factors, i.e., population, researched technologies, wonders, units built and killed, culture, and if spaceship criteria is achieved by the player.
@@ A.2 Mini-game Details
We provide the concrete definition of each mini-game in Table 8. The mini-games belong to three classes -development, battle, and diplomacy, covering a diversity of typical sub-tasks of the Civilization game. The development class of mini-games is closely related to city tile’s combined resource outputs (food, production, and trade), while the battle class of mini-games focuses more on unit movement/attack operations. The player should be aware of the unit attack/defense/HP points to make full use of them. As for the last class of mini-game: diplomacy, it is about the art of trade and negotiation to achieve maximal interests for the player’s civilization.
@@ A.2.1 Victory Conditions, Score/Reward Settings and Difficulty Levels
Based on each mini-game’s characteristics, we design a set of evaluation principles for them. For instance, we assess development mini-games by verifying whether the combined resource outputs (food, production, and trade) from city tiles surpass a predetermined threshold. The success of battle mini-games is determined by the successful annihilation of enemy units or the conquest/defense of designated cities. We determine the achievement of diplomacy mini-games based on whether the agent negotiates favorable agreements with other players, such as exchanging valuable technologies. For the details of reward settings and the computation of victory conditions, please refer to Table 9.
We additionally partition the mini-games into different difficulty levels (easy, normal, and hard) to better evaluate the models’ capability as well as enable curriculum design in future work. Specific criteria for the difficulty level divisions can be found in Table 9.
@@ A.2.2 Mini-game Randomization
To ensure sufficiently diverse and balanced mini-game instances, we conduct a set of randomization processing during mini-game generation. Table 10 gives exact perspectives on which randomization operations are performed for each type of mini-game.
@@
Table 9: Score, victory, reward, and difficulty of mini-games.
@@ B.1 Observation
The tensor environment provides API for tensor models. Observation of tensor environment is a big tensor including all observations of map, unit, city, diplomacy, government, and technology in § A.1.1.
@@ B.2 Action
There are five types of actors in the tensor environment, i.e., unit, city, diplomacy, government, and technology. Each type of actor has its corresponding action space and the overall action space is the joint action space of all actor types. For an actor, each action in its action space is denoted by an action class name associated with parameters. For example, an action in action class Go to a target tile is denoted by the action key: goto_num where goto is the action class name and num is a parameter describing the direction, thus action Go to the north tile is denoted as goto_1 where the parameter 1 corresponds to the north direction.
A chosen action returned to the tensor environment is a triplet including the actor type, actor ID, and the action key. For example, assuming action Go to the north tile is chosen for a unit whose ID is 121, then the action triplet returned to the tensor environment is (unit, 121, goto_1).
@@ C.1 Observation
Please refer to Figure 10 as an example of an actor’s observation dictionary. Zoomed-in observation corresponds to a $5\times 5$ mini-map centered on the actor’s location, thus it contains 25 tiles. Zoomed-out observation corresponds to a $15\times 15$ sub-map. In this example, we define that a block is made up of $5\times 5$ tiles, thus the above $5\times 5$ mini-map is the block where the actor is currently located, and the remaining tiles in zoomed-out observation make up the 8 surrounding blocks. Based on these settings, we get the observation of the actor. Each tile or block has a list containing its status, terrain, and the infrastructures on it. Besides, the information on units and cities, as well as their owners is also on the list. Unit and city owners are also associated with diplomacy tags, for example, units on current_tile belong to myself player; 3 cities in block_north_1 belong to an Alliance player.
@@ C.2 Action
The action returned to the language environment is a triplet including the type of the actor, actor ID, and the name of the chosen action in available_actions. For example, in Figure 10, the type of the actor is unit, its actor ID is 121, assuming that action Go to the north tile has been chosen, then the final action return to language environment is (unit, 121, move North).
@@ Appendix D Estimation of State / Action Space Sizes
On Turn 5. Under ruleset Civ2Civ3, each nation started with 2 Settlers, 2 Workers, and an Explorer. On turn 5, each Settler has acted at most 5 times, so it could appear at 121 possible places as itself, and could build a city at 81 positions. Each city originally had no less than 5 building options, and 24 working options. Each Worker, at the same time, could occur at similar positions, and when some work is done, the range it occurs will be smaller. Explorers could move 3 tiles per turn. So state space for 2 Settlers (or the cities they turned into) is about $\frac{1}{2}((121+81*5*24)^{2}-1521*5*24-121)\approx 5\times 10^{7}$ large (square of single Settlers status, counted in cases: a city is not built / a city is built, with city options, then excluding all conflicting cases where cities are too close to each other. Halving means two Settlers are identical in practice). Workers has states about $\frac{1}{2}((121+81*3)^{2}-121-81*3)=66066$. Explorer contributes $96131^{2}=$ 961 states. So total state space is of size about $3\times 10^{15}$.
@@ E.2.1 Challenges
Sparse and complex observation. In the early stages of the game, there are extensive unseen areas in the player’s field of view. This makes it necessary for early-stage agents to first learn to efficiently explore the environment using powerful prior knowledge, posing a significant challenge in constructing these agents. Additionally, the vast map size (4000-tile maps for our full game setup) and complex information on each tile overwhelm the context window of all language models. Thus, LLM agents have to prioritize reading the map properly and efficiently.
@@ E.2.2 Structure
We adopted an AutoGPT-like approach to build a baseline language-based agent. AutoGPT [72] is a memory-equipped automated LLM agent capable of making autonomous decisions in context. The following sections will delve into observation, reasoning, commands, and structure separately. ion.